Please help me answer all 3 questions in this case. The numbers and questions asked, in this case, is different than the other cases similar to it, so please don't copy and past other answers from the same case. Also if you could please provide steps to how you're getting the answers, if possible show it in excel with formulas, please. And also please mark the TOTAL COST OF THE MOVE FOR QUESTION 1.

CASE 4-2 Hardee Transportation (B) One of Jim OBriens customers has presented him with an opportunity for a signifi- cant amount of freight moving into a new market for Hardee. Hardee is a truckload carrier primarily moving freight in the East/West market in the United States. Although it has some movements in and out of Canada and Mexico, Hardee has focused on moving freight in eastward and westward directions. Hardee has dispatch centers located throughout the United States which have some dock capacity The new move would be between Pittsburgh and Miami. Hardee has avoided this market because of the lack of backhaul opportunities that exist outbound from Florida. However, this new move offers a significant increase in volume for Hardee. A compli- cating factor in this move is the request that Hardee perform sorting and segregation at its dispatch centers. Each shipment will consist of straight (one product) pallet loads of various types of consumer goods freight destined for a retailers distribution center in Miami. Sorting and segregation at Hardees locations would consist of breaking the pallets and sorting the freight by the retailers store locations, then repalletizing into rainbow (mixed products) pallets for each store. Hardee has never experienced this type of request before. Jim knows that he needs to put some type of costs to this move to make sure that the moves are profitable. Because of the large volume involved, not covering Hardees costs in pricing could result in large losses for Hardee. The relevant information for costing this move is as follows: Equipment Cost Data Equipment Purchase Price 1. Line-haul tractors $80,000 2, Line-haul trailers $24,000 L. Tractors 5-year straight line Trailers- 8-year straight line Interest 1. Tractors-10 percent APR for 5 years 2. Trailers 10 percent APR for 8 years Fuel

1. $2.10 per gallon for diesel 2. Line-haul tractors 6.5 miles per gallon Labor 1, Line-haul drivers $0.42 per mile 2. Pick-up and delivery (PUD) drivers $30 (fully loaded) per hour 3. Dock workers $25 (fully loaded) per hour

Miscellaneous I. Insurance cost-$0.05 per mile . Maintenance cost- $0.15 per mile . Billing cost $5.00 per freight bill 4. Administrative/overhead cost- 8 percent of total cost of move 5. Dock facility cost- $15 per hour 6. Line-haul vehicle averages 45 mph between origin and destination. Route and Time of Move The shipment (40,000 pounds) originates at a customer location in Pittsburgh, located 20 miles from Hardees dispatch center. A PUD driver is dispatched from the Hardee location at 8:30 a.m. on January 12, 2010, and arrives at destination at 9:00 a.m the same day. The shipment is loaded from 9:00 a.m. to 12:00 p.m. The PUD driver departs the customer location at 12:00 p.m. and arrives back at the Hardee dispatch center at 12:30 p.m The sort process starts at 12:30 p.m. and ends at 8:30 p.m. on January 12. It requires unloading the trailer, sorting, and repalletizing the load. This operation requires two dock workers, each working the same trailer for 8 hours in the dispatch center The line-haul portion begins with the vehicle being dispatched from the Pittsburgh location at 8:30 p.m. on January 12 and traveling to Charlotte, North Carolina, a dis- tance of 481 miles, and arriving at Charlotte at 7:12 a.m. on January 13. The driver rests from 7:12 a.m. until 3:12 p.m. The trip continues with the vehicle departing Charlotte at 3:12 p.m. on January 13 and traveling to Jacksonville, Florida, a distance of 399 miles, arriving at Jacksonville at 12:06 a.m. on January 14. The driver rests from 12:06 a.m. until 10:06 a.m. The line-haul portion concludes with the vehicle departing Jacksonville at 10:06 a.m. and traveling to the customers location in Miami, a distance of 369 miles, and arriving at the distribution center at 6:18 p.m. on January 14 The line-haul driver stays with the vehicle while it is being unloaded (2 hours un load time). The driver then deadheads at 8:18 p.m. from the customers distribution center and arrives at a Hardee dispatch center located in Miami at 8:48 p.m., a distance of 15 miles from the distribution center

Hourly Principal and interest Calculator Loan Amount Interest Rate Days in a Year Tractor $80,000.00 1096 365 Trailer $24,000.00 1096 365 Length of Loan (years) Monthly Payment ($1,699.76) PMT(C9/12,C12*12,C8 ($364.18) PMT(E9/12,E12*12,E8) $4,370.16) E13 12 Annual Payment Hourly Payment $20,397.16) $2.33 C13 12 C16/(24 C10) $0.50 E16/(24 E10)
Answer #1

TOTAL COST= $1,243.27
PER cwt= $3.45
PER Revenue Mile= $1,243.27/1295 miles = $0.96


Hardee Costs
Pick up costs:
Labor: 4 hrs x $30/hr = $120
Fuel: 40 miles / 6.5 mpg x $2.10/gal. = $12.92
Insurance & Maintenance: 40 m. x $0.20/m = $8.00
Tractor Depreciation: 4hrs. x $1.83/hr = $7.32
Trailer Depreciation: 4 hrs. x $0.34 = $1.36
Billing cost:$5
Total Pick Up Costs: $154.50


Sorting Costs:
Labor: 16hrs x $25/hr = $400
Tractor Depreciation: 8hrs x $1.83/hr = $14.64
Trailer Depreciation: 8hrs x $0.34 = $2.72
Total Sorting Costs: $417.36
Line-Haul cost
Depreciation: tractor: 46 hr @$1.83/hr = $84.18
Trailer: 46 hr @$0.34/hr = $15.64
Interest: Tractor: 46 hr@$2.33/hr = $107.18 Trailer: 46hr @ $0.50= $23.00
Fuel: 1249 @ $0.32/mile =$399.68
Labor: 1249 @ $0.42/mile = $54.18
Maintenance: 1249@ $0.15 = $187.35
Insurance: 1249@ $0.03 = $37.47
Total Line-Haul Cost = $908.68

Delivery Costs:
Tractor:2.30hr @ $1.83/hr = $4.21 Trailer:2.30hr @ $0.34/hr= $0.78
Interest: Tractor: 2.30hr @ $2.33/hr=$5.36 Trailer: 2.30hr @$0.50/hr= $1.15
Fuel: 15 miles@ $0.32/mile= $4.80
Labor: 2.30hr @ $30.00/hr= $69.00
Maintenance:15 miles @ $0.15/mile= $2.25
Insurance: 15 miles @ $0.03/mile= $0.45
TOTAL Delivery Cost= $88.00